• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ellie

    Log entry time 16:04:56 on October 06, 2009

    Entry number 293497

    keyword=Day Shift Summary

    10/06/09 Day Shift Summary

    On shift:
    SL: Ellie Long
    TO: Mustafa Canan
    3rd: Arun Saha

    · 08:00 Good morning, Hall A! Starting shift with problems in Left Q1, Right Q1 and Right Dipole. We are in controlled access so that the techs can look at the spectrometers. IHWP=OUT, Slug #19

    · 11:45 Techs have finished working on the HRS magnet repairs. Both left and right are ramping up.

    · 11:41 Having problems with Right Arm LeCroy HV. Restarting HVGUI didn't fix the problem. Called Brad who is going to go down in the hall since we still have controlled access to take a look at it.

    · 11:45 Controlled access continues for PVDIS work and fixing of LeCroy HV.

    · 12:50 Work in hall is done, going back to Beam Permit and doing Beam Recovery Procedure.

    · 13:15 We are back up and running production (Note: RHRS is still ramping up, we're only taking production with the left arm). Spot++ at 100uA looks good.

    · 15:35 The RHRS Dipole is back on (P0=3.14177)

    · 16:00 Shift ends in Production Mode with run 13012 still going.

    Production Runs:
     Run #  Beam Current  # of Good Pairs  Comments 
    13009100uA19.4kRHRS Dipole ramping up
    13010100uA17.4kRHRS Dipole ramping up
    13011100uA6.0kRHRS Dipole ramping up
    13012100uAStill Running as shift endsRHRS Dipole is back up (P0=3.14177)